### v1

@@ 5.2 Estimating the Braking Parameters
The specification of the braking force in Eq. (12) is incomplete without estimates for the braking coefficient $\gamma$ and threshold ${a}_{\text{\emph{c,min}}}$. The coefficient $\gamma$ can be determined using experimental data reported in Sandbakk et al. [22], who studied elite female skiers applying both skid and step turn techniques while navigating curved downhill sections of a ski course. These authors measured the change in mechanical energy ($\Delta e$) between the start/end of a turn, average skier speed ($\overline{v}$), and “trajectory” or distance travelled ($D$). Their data are summarized in the first four rows of Table 3 for four different combinations of turn geometry ($R$, $D$) and skid/step technique.
Table 3: Measured data from [22] on elite female skiers (with average mass $m=60$ kg) performing several different skid/step turns. The first 4 rows list the original data while the last row contains our estimate of the braking coefficient $\gamma$ based on Eq. (20). Skiers employed the skid technique in the two tighter turns ($R=9,12$), and the step technique in the wider turns ($R=12,15$).
@@
$\Delta e$ kg  m2/s2 26.2 25.4 14.9 17.2
$\gamma$ 0.0060 0.0055 0.0024 0.0028
$\underbrace{\text{\hskip 28.45274pt}}$${\text{\hskip 28.45274pt}}$
${\gamma}_{\text{\emph{skid}}}\approx 0.0058$ ${\gamma}_{\text{\emph{step}}}\approx 0.0026$
Assuming that the braking force is constant throughout each turn, the mechanical energy (or work) can be estimated using the basic physical principle “work equals force times distance”:
$\underbrace{~{}\Delta e{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\big{|}}}_{\text{work}}\;=\;\underbrace{\big{(}m\gamma\kappa\overline{v}^{2}\big{)}}_{\text{force}}\;\cdot\underbrace{~{}D{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\big{|}}}_{\text{distance}}\qquad\Longrightarrow\qquad\gamma=\frac{R\,\Delta e}{m\overline{v}^{2}D}\raisebox{2.84544pt}{.}$ (20)
Substituting the average skier mass $m=60$ kg and other parameters from Table 3 into Eq. (20) yields $\gamma$ values reported in the final row of the table. Turning technique clearly has a significant effect on the braking coefficient, with $\gamma$ for the step turn equal to roughly half that for the skid turn. On the other hand, $\gamma$ depends only very weakly on the turn radius $R$ and length $D$, so we choose the average of the two braking coefficient values ${\gamma}_{\text{\emph{skid}}}=0.0058$ and ${\gamma}_{\text{\emph{step}}}=0.0026$ for the skid- and step-turn techniques, respectively.
The other parameter in our braking model is the threshold ${a}_{\text{\emph{c,min}}}$, which cannot be determined from [22] since they did not consider the different situations when braking turn techniques could be applied. However, the Homologation Manual provides a list of maximum limits for centripetal acceleration [8, p. 26] that are permitted within the various Nordic ski race formats. The smallest two limits allowed in all races are $a_{c}=2$ and 5 m/s2, which suggests two possible values for ${a}_{\text{\emph{c,min}}}$ that will be investigated in the next section.
@@ 6 Example 4: 3D Simulations of the Ole Course
We are now ready to simulate a skier moving along an actual 3D ski course. We return to Example 3 and the 4.2 km Ole course from Toblach, and solve the ODE system (14)–(15) by making use of all Hermite spline computed by setup3d (not just those for $Z$, $S$ and $\Theta$). Six simulations are performed taking values of the braking coefficient $\gamma=0$ (no braking), $0.0026$ (step turn), and $0.0058$ (skid turn), each being applied with two threshold values ${a}_{\text{\emph{c,min}}}=2,5$. Otherwise we use the same MSH baseline parameters from Table 1 that were used in 2D simulations.
@@
The corresponding completion times are summarized in the first two rows of Table 4 (labelled “Ole”) where it is clear that using either the skid or step turn increases the finish time by a small margin, on the order of one second or less. The slowest time results from applying the skid turn with threshold ${a}_{\text{\emph{c,min}}}=2$ (requiring an extra 1.12 s) whereas the skier is slightly faster when applying a larger threshold of ${a}_{\text{\emph{c,min}}}=5$. Note that the total time required without braking matches the 2D result of 819.9 s reported in Section 4.3, because the 2D and 3D ODE systems are identical when $\gamma=0$. For comparison purposes, we repeat these same six simulations using GPS data from another 3.9 km FIS-rated course from Toblach called “Stephanie,” which features more tight curves than the Ole course. The corresponding times are given in the last two rows of Table 4, which are also only slightly slower than the “no brake” case; however, the differences are slightly larger than the Ole course since there are more tight curves on which braking forces are applied.
@@
Table 4: Simulated race times for the Ole and Stephanie courses in Toblach. The results are presented for three choices of braking technique (no braking, step turn, skid turn) and two braking thresholds (${a}_{\text{\emph{c,min}}}=2,5$). The figures in parentheses indicate the corresponding changes relative to the time without braking.
Course ${a}_{\text{\emph{c,min}}}$ No brake Step turn Skid turn
Ole (4.2 km) 2 819.94 820.29 ($+$0.35) 820.71 ($+$0.77)
5 820.14 ($+$0.2) 820.39 ($+$0.45)
Stephanie (3.9 km) 2 801.57 802.60 ($+$1.03) 803.92 ($+$2.35)
5 802.45 ($+$0.88) 803.58 ($+$2.01)
@@
Returning to the Ole results, Fig. 9 provides a more detailed look at the effects of braking by plotting the difference in projected distance, $\xi_{2D}-\xi_{3D}$, between the case with no braking ($\xi_{2D}$) and that with skid/step turns ($\xi_{3D}$). The tightest downhill curves occuring in the final third of the course can be identified with the largest discrepancies between the three results. At the end of the course, the skier using skid turns is 7 m behind the idealized athlete who applies no braking, whereas using step turns leaves the skier only 3 m behind. This emphasizes how a simple adjustment in braking technique on only a few difficult turns could allow one skier to make a significant gain over another. As mentioned earlier in Section 5.1 (in reference to Bjørgen’s 2010 gold medal performance), such small differences can be of critical importance in a race scenario where competitors are skiing neck-and-neck and must jockey for any possible advantage.
Refer to captionABC
@@
Refer to caption (b) Solution components versus time (c) Zoomed curvature in Region B Refer to captionABC Refer to caption
Figure 10: (a, Top) The braking force is displayed along the scaled course profile in a 3D view of the solution with $\gamma={\gamma}_{\text{\emph{skid}}}$ and ${a}_{\text{\emph{c,min}}}=$2. The skier brakes much more often during the second half, where the course is mostly downhill. (b, Bottom Left) Plots of $F_{b}$, $v$, $\kappa$ and $\theta$ versus time demonstrate how the braking force arises from a combination of skier speed and track curvature. (c, Bottom Right) Zoomed-in view of the curvature plot in region B.
  • Region A: This starts with a tight curve on an uphill climb at $t\approx 50$ s, followed by a steep downhill that also ends on a highly curved section. Here, the skier only brakes near the bottom of the hill because they are not travelling fast enough on the initial uphill to exceed the braking threshold.
  • Region B: This stretch contains the steepest downhill which also has two tight curves near the hill crest and base, coinciding with the two spikes in $\kappa$ near $t\approx 572$ and $595$ s. The first curve occurs just over the crest when the skier is only beginning to speed up but curvature is large enough to exceed the braking threshold and generate a very small spike in the braking force. A similarly sized “blip” in $F_{b}$ occurs at $t\approx 590$, after the course has flattened but the skier is still travelling at very high speed. This is still before the reaching the tight curve at $t\approx 595$, but there is an extended section beforehand with small-to-moderate curvature that is sufficient to also pass the braking threshold (see Fig. 10c). This example demonstrates how large speed (or curvature) on its own can generate a small braking force.
@@
By adding curvature into our model, we have shown how to quantify the effects of applying various braking techniques. When looking only at the effects on the total race time, the braking force may seem almost negligible on courses like Ole, but our study shows that choosing a more effective turning technique in a few strategic locations can still provide a gain on the order of a few meters. Additionally on more convoluted courses like Stephanie, this impact could be even greater. This simple addition to the model provides opportunities for collaboration with coaches, athletes, and sports scientists to test our model predictions against real athlete data, and to investigate different race strategies with a view to optimizing race results.
Exercise 16.
@@
  • They apply the “snowplow” technique on all downhills regardless of speed or curvature, which has a much larger braking parameter of roughly $\gamma=10{\gamma}_{\text{\emph{skid}}}$.
Adjust the parameters for a skier on the Ole course, and aim for a finish time that is twice the baseline result.
Exercise 17.
Download GPS data for the 1.4 km “Albert Sprint” course (also called “Stadium Track”) from the Dolomiti NordicSki website [6] and simulate a skier in 3D with curvature and braking. Compare your results to the 2D simulation from Exercise 13.
@@ 7 Discussion
  • Data analysis and cleaning: to investigate and correct the random noise and other measurement errors that are inherent in real input data.
@@
Finally, we should emphasize that this work is still only a preliminary step towards developing a truly realistic model for a very complex and multi-faceted sport, and there are a host of opportunities for further work. One specific area is the study of athlete fatigue in medium- to long-distance events, and how to incorporate fatigue into the power model (beyond the naïve approach we suggested in Example 12). Another natural extension is to extend our simplistic approach for handling braking turns to incorporate “look-ahead,” where a skier anticipates arriving at a tight curve by initiating their braking technique well in advance. There are many other fascinating aspects of Nordic skiing that we have so far ignored, such as the metabolic processes behind muscle power generation, or the variability in course conditions due to weather-induced snow conditions, or track grooming, or athlete race strategy, to name just a few. In terms of strategic decision-making, mathematics is especially well-suited to drive advances in understanding how specific choices of technique or training regimen could optimize race performance under a variety of parameter regimes. Existing work on applying methods from optimization to maximize performance in running races [1, 11] could be extended naturally to Nordic skiing along similar 4lines to what was already done in [7].
@@ Acknowledgements
We thank Kayden Sim (GIS Specialist, Forsite Consultants) and Jake Weaver (Head Coach, Hollyburn Cross-Country Ski Club, North Vancouver, BC). We also appreciate the advice of FIS Technical Delegates and Course Designers John Aalberg (Sooke, BC) and Al Maddox (Atikokan, ON), who assisted us with details regarding course homologation.
@@
  • S1
    Exercise solutions (suppl1Solutions.pdf, 10.3 Mb)
  • S2
    Summary of Matlab codes and data files (suppl2CodesAndData.pdf, 140 Kb)
  • S3
    Matlab codes (suppl3Matlab.zip, 13 Kb)
  • S4
    Data files (suppl4DataFiles.zip, 12 Kb)
  • S5

### v2

@@ 5.2 Estimating the Braking Parameters
The specification of the braking force in Eq. (12) is incomplete without estimates for the braking coefficient $\gamma$ and threshold ${a}_{\text{\emph{c,min}}}$. The coefficient $\gamma$ can be determined using experimental data reported in Sandbakk et al. [28], who studied elite female skiers applying both skid and step turn techniques while navigating curved downhill sections of a ski course. These authors measured the “trajectory” ($D$, or distance travelled), average skier speed ($\overline{v}$), and the change in specific mechanical energy ($\Delta e$, or energy per unit mass) between the start/end of a turn. Their data are summarized in the first four rows of Table 3 for four different combinations of turn geometry ($R$, $D$) and skid/step technique.
Table 3: Measured data from [28] on elite female skiers (with average mass $m=60$ kg) performing several different skid/step turns. The first 4 rows list the original data while the last row contains our estimate of the braking coefficient $\gamma$ based on Eq. (20). Skiers employed the skid technique in the two tighter turns ($R=9,12$) and the step technique in wider turns ($R=12,15$).
@@
$\Delta e$ J/kg 26.2 25.4 14.9 17.2
$\Delta E=m\Delta e$ J 1573 1525 892 1033
$\gamma$ 0.3582 0.3324 0.1460 0.1686
[…]
${\gamma}_{\text{\emph{skid}}}\approx 0.3453$ ${\gamma}_{\text{\emph{step}}}\approx 0.1573$
Assuming that the braking force is constant throughout each turn, the mechanical energy (or work) can be estimated using the physical principle that “work equals force times distance”:
$\underbrace{~\Delta E{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\big|}}_{\text{work}}\;=\;\Delta e\cdot m\;=\;\underbrace{\big(m\gamma\kappa\overline{v}^{2}\big)}_{\text{force}}\;\cdot\underbrace{~D{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\big|}}_{\text{distance}}\qquad\Longrightarrow\qquad\gamma=\frac{\Delta e}{\kappa\overline{v}^{2}D}=\frac{R\,\Delta e}{\overline{v}^{2}D}\raisebox{2.84544pt}{.}$ (20)
Substituting the average skier mass $m=60$ kg and other parameters from Table 3 into Eq. (20) yields $\gamma$ values reported in the final row of the table. Turning technique clearly has a significant effect on the braking coefficient, with $\gamma$ for the step turn equal to roughly half that for the skid turn. On the other hand, $\gamma$ depends only very weakly on the turn radius $R$ and length $D$, so we choose the average of the two braking coefficient values ${\gamma}_{\text{\emph{skid}}}=0.3453$ and ${\gamma}_{\text{\emph{step}}}=0.1573$ for the skid- and step-turn techniques, respectively.
The other parameter in our braking model is the threshold ${a}_{\text{\emph{c,min}}}$, which cannot be determined from [28] since they did not consider the different situations when braking turn techniques could be applied. However, the Homologation Manual provides a list of maximum limits for centripetal acceleration [10, p. 26] that are permitted within the various Nordic ski race formats. The smallest two limits allowed in all races are $a_{c}=2$ and 5 m/s2, which suggests two possible values for ${a}_{\text{\emph{c,min}}}$ that will be investigated in the next section.
@@ 6 Example 4: 3D Simulations of the Ole Course
We are now ready to simulate a skier moving along an actual 3D ski course. We return to Example 3 and the 4.2 km Ole course from Toblach, and solve the ODE system (14)–(15) by making use of all Hermite spline computed by setup3d (not just those for $Z$, $S$ and $\Theta$). Six simulations are performed taking values of the braking coefficient $\gamma=0$ (no braking), $0.3453$ (skid turn), and $0.1573$ (step turn), each being applied with two threshold values ${a}_{\text{\emph{c,min}}}=2,5$. Otherwise we use the same MSH baseline parameters from Table 1 that were used in 2D simulations.
@@
The corresponding completion times are summarized in the first two rows of Table 4 (labelled “Ole”) where either the skid or step turn increases the finish time by roughly 10–30 s, which is a significant difference in a race scenario. The slowest time results from applying the skid turn with threshold ${a}_{\text{\emph{c,min}}}=2$ (requiring an extra 29.6 s), whereas increasing the threshold to ${a}_{\text{\emph{c,min}}}=5$ (reducing the frequency of braking) allows the skier to finish slightly faster. Note that the total time required without braking matches the 2D result of 819.9 s reported in Section 4.3, because the 2D and 3D ODE systems are identical when $\gamma=0$. For comparison purposes, we repeat these same six simulations using GPS data from another 3.9 km FIS-rated course from Toblach called “Stephanie,” which features more tight curves than the Ole course. The corresponding times are given in the last two rows of Table 4, which exhibit time differences from the “no brake” case that are more than double that from the Ole course. This result is not surprising since the Stephanie course has more tight curves on which braking forces are applied.
@@
Table 4: Simulated race times for the Ole and Stephanie courses in Toblach. The results are presented for three choices of braking technique (no braking, skid turn, step turn) and two braking thresholds (${a}_{\text{\emph{c,min}}}=2,5$). The figures in parentheses indicate the corresponding changes relative to the time without braking.
Course ${a}_{\text{\emph{c,min}}}$ No brake Skid turn Step turn
Ole (4.2 km) 2 819.9 849.5 (++29.6) 837.5 (++17.6)
5 837.6 (++17.7) 830.3 (++10.4)
Stephanie (3.9 km) 2 801.6 864.5 (++62.9) 840.8 (++39.2)
5 852.4 (++50.8) 834.3 (++32.7)
@@
Returning to the Ole results, Fig. 9 provides a more detailed look at the effects of braking by plotting the difference in projected distance, $\xi_{2D}-\xi_{3D}$, between the case with no braking ($\xi_{2D}$) and that with skid/step turns ($\xi_{3D}$). The tightest downhill curves that occur in the final third of the course can be identified with the largest discrepancies between the three results. At the end of the course, the skier using skid turns finishes roughly 240 m behind the idealized athlete who applies no braking, whereas a skier using the more efficient step turns lags by only 120 m. This emphasizes how a simple adjustment in braking technique on only a few difficult turns could allow one skier to make significant gains over another. As mentioned earlier in Section 5.1 (in reference to Bjørgen’s 2010 gold medal performance), even very small differences can be of critical importance in a race scenario where competitors are skiing neck-and-neck and must jockey for any possible advantage.
@@
[…]
Figure 10: (a, Top) The braking force is displayed along the scaled Ole course profile in a 3D view of the solution with $\gamma={\gamma}_{\text{\emph{skid}}}$ and ${a}_{\text{\emph{c,min}}}=$2. The skier brakes much more often during the second half, where the course is mostly downhill. The sections marked A,B,C indicate tightly curved downhill sections where braking forces are applied. (b, Bottom Left) Plots of $F_{b}$, $v$, $\kappa$ and $\theta$ versus time demonstrate how the braking force arises from a combination of skier speed and track curvature. (c, Bottom Right) Zoomed-in view of the curvature plot in region B.
  • Region A: This starts with a tight curve on an uphill climb at $t\approx 50$ s, followed by a steep downhill that also ends on a highly curved section. Here, the skier only brakes near the bottom of the hill because they are not travelling fast enough on the first short downhill section to exceed the braking threshold.
  • Region B: This stretch contains the steepest downhill section on the course which also has two moderately tight curves near the hill crest and base, corresponding to the two prominent spikes in $\kappa$ near $t\approx 580$ and $602$ s. The first curve is encountered just over the hill crest when the skier is only beginning to speed up but their centripetal acceleration is still sufficient to exceed the braking threshold and generate a moderate braking force with $F_{b}\approx 1$. In contrast, the similar curvature near the base of the hill occurs once the course has flattened and the skier has slowed down enough that no braking is triggered. Nevertheless, there is a time in between (near $t\approx 597$ s) when the skier encounters only a modest curve on the steep downhill (circled in Fig. 10c), yet is moving sufficiently fast to induce a similarly-sized spike in $F_{b}$. This example illustrates how braking forces can be generated by a very fast moving skier on a moderate curve, as well as on very tight curves at moderate speeds.
@@
By adding curvature and a simple braking force into our model, we have shown how to quantify the effects of applying various braking techniques and that braking can have a significant impact on the finishing time. Additionally on more convoluted courses like Stephanie, this impact could be even greater. This simple addition to the model provides opportunities for collaboration with coaches, athletes, and sports scientists to test our model predictions against real athlete data, and to investigate different race strategies with a view to optimizing race results.
Exercise 17.
@@
  • They use the “snowplow” technique on all downhills regardless of speed or curvature, employing the much larger braking parameter $\gamma=10{\gamma}_{\text{\emph{step}}}$.
For the Ole course, adjust the power parameter to obtain a finish time that is twice the baseline result.
Exercise 18.
Download GPS data for the 1.4 km “Albert Sprint” course from the Dolomiti NordicSki website [8] and simulate a skier in 3D with curvature and braking. Compare your results with those from the corresponding 2D simulation in Exercise 14.
Exercise 19.
Consider an artificial 3D course that is built using the function $r(\theta)=100(1+\frac{1}{2}\cos n\theta)$, which describes a curve in the $x,y$-plane in polar coordinates (for $n\geqslant 1$ an integer). To introduce changes in elevation, this curve is then rotated an angle of $10^{\circ}$ about the $y$-axis. This yields a smooth “star-shaped” path having $n$ rounded points or lobes that become more highly curved as $n$ increases, and with gradients varying between $\pm 10^{\circ}$. The Supplementary Information includes a MATLAB function called starcourse.m that generates $(x,y,z)$ coordinates using:
[x, y, z] = starcourse(3, 100)
where the first argument is the number of lobes (here, $n=3$) and the second argument indicates that 100 data points should be used to define the course. Modify the code
setup3d.m to use this command to define the course rather than reading a GPS data file, and then run the code for a selection of $n$ values between 1 and 20. The code will generate some anomalous or erroneous results for different values of $n$. Discuss the results, explain what is going wrong when errors arise, and then try to correct the code.
@@ 7 Discussion
  • Data analysis and cleaning: to investigate and correct for missing data points, highly non-uniform point spacing, and other measurement errors that are unavoidable with real data sources.
@@
In closing, we should emphasize that this work is still only a preliminary step towards developing a truly realistic model for a very complex and multi-faceted sport, and there are a host of opportunities for further work. One specific area is the study of athlete fatigue in medium- to long- distance events, and how to incorporate experimental observations on positive pacing into the power model beyond the naïve approach suggested in Exercise 13. Another natural extension is to generalize our simplistic approach for handling braking turns to incorporate “look-ahead,” where a skier anticipates arriving at a tight curve by initiating their braking technique well in advance. There are many other fascinating aspects of Nordic skiing that we have so far ignored, such as the metabolic processes behind muscle power generation, or the variability in course conditions due to weather-induced snow conditions, or track grooming, or athlete race strategy, to name just a few. In terms of strategic decision-making, mathematics is especially well-suited to drive advances in understanding how specific choices of technique or training regimen could optimize race performance under a variety of parameter regimes. Finally, existing models of athlete dynamics in other competitive sports could provide interesting avenues to explore in the context of Nordic skiing. For example, optimization methods developed to maximize performance in a running race [1, 13] could be extended to skier dynamics along similar lines to what was already done in [9]. Furthermore, in the sport of road cycling athletes propel themselves at much higher speeds where aerodynamic forces dominate [25], which would provide an opportunity to draw interesting comparisons with skiing.
@@ Acknowledgements
We thank Kayden Sim (GIS Specialist, Forsite Consultants) and Jake Weaver (Head Coach, Hollyburn Cross-Country Ski Club, North Vancouver, BC). We also appreciate the advice of FIS Technical Delegates John Aalberg (Sooke, BC) and Al Maddox (Atikokan, ON) who assisted us with details regarding course homologation. Finally, we are very grateful to the two anonymous reviewers for their extensive and insightful comments that allowed us to make significant improvements to our final paper, and provided the inspiration for two additional exercises.
@@
  • S1.
    Exercise solutions (suppl1Solutions.pdf, 11.9 Mb)
  • S2.
    Summary of MATLAB codes and data files (suppl2CodesAndData.pdf, 131 Kb)
  • S3.
    MATLAB codes (suppl3Matlab.zip, 15 Kb)
  • S4.
    Data files (suppl4DataFiles.zip, 13 Kb)
  • S5.
@@
  • S7.
    Movie of 3D simulation for “star-shaped” course in Exercise 19 (suppl7Movie3dStar7.mp4, 2.0 Mb)